Find the value of (x-a)(x-b)(x-c)...(x-z)?
Pls help me out with detailed reasoning and step by step solution?

Grade:11

5 Answers

Arun
5 years ago
Dear student
 
I think that your question is incomplete because there is no information. Please check and repost it. You can attach an image too.
I will be happy to help you.
Surya
5 years ago
i think we can do it like this,

We have to find,

(X-A)(X-B)(X-C)....(X-Y)(X-Z)

= (X-A)(X-B)(X-C)(X-D)....(X-W)(X-X)(X-Y)(X-Z)

= (X-A)(X-B)(X-C)(X-D)....(X-W)(0)(X-Y)(X-Z) (since X-X = 0)

= 0 [Product of any term multiplied with zero always results in zero].

Hence the answer is zero.

Arun
5 years ago
Dear Surya
 
I thought that 1st term i.e. x is a variable while a,b, c,....z are constant relative to that x.
 
So I thought X- x can't be zero.
 
Otherwise if you mean both x are same then you are correct.
Kushagra Madhukar
3 years ago
Dear student,
 
 Since (x – x) is one of the term as per the information, and both x mmeans the same thing hence this term is = 0
Hence, the final answer = 0
 
Thanks and regards,
Kushagra
amit yadav
2 years ago

In the sequence of multiplication there is term (x-x) = 0

Anything multiplied by zero is always Zero.

So, the value of whole term is 0.